News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 30 March 2001 6.30pm EST

During the last week, Chandra completed the observing schedule as planned.

The Chandra team were kept well informed by JPL about several work stoppages at the Canberra DSN site. Chandra operations were not impacted however we will continue to monitor the situation via JPL until resolved.

A class X solar flare was detected on 3/29 resulting in a series of coronal mass ejections (CMEs), however neither the EPHIN threshold or the orbital fluence threshold were exceeded and the loads were not halted due to this event. The solar levels continue to be closely monitored.

Of note last week was a visit by the Chandra Operations leads to meet with the GSFC Hubble and Institute Operations team. The purpose of the visit was to share operations approaches with topics including anomaly response, procedure development, CM, mission planning and ground system. The visit was very valuable and a visit by the Hubble team to the OCC will be planned.

Of note next week will be a live television broadcast from the OCC at 1pm EDT on April 3 on NASA-TV and participating public TV stations. The "Passport to Knowledge Program" is aimed at school children and will include a live Q&A via email between with Chandra astronomers and students.

The planned observations for the coming week are shown below and include an observation of the Crab Nebula coordinated with the Hubble Space Telescope. Real-time procedures are planned for April 6 and 7 to insert and retract the LETG.

NGC 4314                ACIS-S       Apr 2
NGC4111                 ACIS-S
1928+738                ACIS-S       Apr 3  
1641+399                ACIS-S
XMM 13hr survey field   ACIS-I       Apr 4
Radiation Belt
V 1121+2327             ACIS-I       Apr 5  
Crab Nebula             ACIS-S       Apr 6  
PKS2155-304             HRC-S/LETG
4U 0241+61              ACIS-S/LETG
Radiation Belt                       Apr 7  
0836+299                ACIS-S
ISO NW #1               ACIS-I       Apr 8


All spacecraft subsystems continued to operate nominally.

Return to Status Reports